OpenStudy (anonymous):

Multiply. Assume that all expressions are defined. x/15∙x^7/2x∙20/x^4

OpenStudy (anonymous):

Step 1: \[\frac{ x(x^{7})(20) }{ 15(2x)(x^{4}) }\] Multiply it all the top of the fractions (called numerator) together. Multiply all the bottom (denominator) together. Tell me what you get

OpenStudy (anonymous):

im so lost idk i need alot of help

OpenStudy (anonymous):

what is x times x^7

OpenStudy (anonymous):

? 2x^7

OpenStudy (anonymous):

Think of x as x^1 x^1 times x^7 = x^ (1+7) = x^8 :) Let's try another one outside of this problem, to make sure you get the concept? Outsite info: What's x^ 2 times x^4?

OpenStudy (anonymous):

that is not one of my choices

OpenStudy (anonymous):

this is part of the problem only, we haven't reached the answer yet.

OpenStudy (anonymous):

back to the problem: what is 20 times x^8 not answer yet